a. Is distillation a purification or a characterization technique? Explain.

b. Where the thermometer bulb should be placed in order to measure the temperature of the hot vapor accurately during distillation?

c. What is the purpose of the “packing” in a fractional distillation column?

d. Why should a distillation flask never be more than ¾ full before starting a distillation?

e. You are fractionally distilling 100 mL of a 1:1 solution of cyclohexane and toluene. After 2 mL has distilled you collect a sample for analysis. It is determined that your sample contains a mixture that is 92:8 cyclohexane/toluene. Calculate the number of theoretical plates if the relative volatility (α) of the original solution is 2.33.

f. From the results of your graph, discuss which method of distillation would be preferred for separating acetone and ethanol and why
Answer #1

(a)

Distillation: This is purification method of mixture liquid which based on boiling point of the liquid.There is five type of distillation as follows:

Simple distillation, Fractional distillation, Vacuum distillation, Steam distillation and separating funnel.

So, distillation is purification technique of mixture of liquid.

(b)

In distillation, the thermometer uses to measure the temperature of the vapor that about to be condensed.So, put the thermometer such a way that entire mercury reservoir is immersed in the rising vapors and that measure accurate temperature.

(c)

The column of fractional distillation is packed with glass beads to provide a maximum possible surface area for vapor to condense on.

(d)

We used to avoid fill a flask completely, There is a chance for the liquid to come out from the flask without vaporizing contaminant the distillate.

(e)

calculate relative volatility\alpha =Y(1-X)/X(1-Y)

where X and Y are mole fraction of cyclohexane and toluene.

the mole fraction of cyclohexane=92/92+8=0.92

the mole fraction of toluene=8/92+8=0.08

\alpha =0.08(1-0.92)/0.92(1-0.08)=0.0076

numbers of plate=relative volatility/calculated volatility=2.33/0.0076=308

(f)

the boiling point of ethanol is 78.37 0C and acetone is 560C. there is a difference in their boiling point.so we can use fractional distillation technique. but ethanol contains lots of water molecule we add some amount of MgSO4, it dries the ethanol solution and goes for fractional distillation.It is a separation technique for a mixture of liquids on the basis of their boiling point. which solvent has the lowest boiling point, will vaporize first.